# Paying to Do Better: Games with Payments between Learning Agents

### v1

@@ Appendix B Prisoner’s Dilemma Games
There are various ways to show this. Consider the following induction argument. The game has at least one order of iterated elimination of strictly dominated strategies. Index the players according to one such order, such that player $1$ has a dominant strategy, player $2$ has a dominant strategy in the sub-game where player $1$ plays his dominant strategy, and so on. Additionally, index the corresponding dominant strategies of the players using the same order, so that $a_{1}$ is the first action that is taken, $a_{2}$ is the second, and so on. Clearly, in any CCE, player $1$ plays his dominant strategy $a_{1}$; otherwise he has positive regret. Our induction step is simple: Fix any CCE. Given that players $11,...,k-$ 1 play actions $a_{1},...,a_{k-1}$, player $k$ has a strict best response to play $a_{k}$. By induction, this is true up to the last player $n$,so we have that our arbitrary CCE is the pure Nash equilibrium profile $(a_{1},...,a_{n})$. Therefore, regret-minimization dynamics converge to it with high probability. ∎
@@
To approach the outcome described in the introduction as an equilibrium, we need to adjust the payment policy of player $1$ as follows. Player $1$’s agent pays player $2$’s agent the maximum payment whenever player $1$’s agent cooperates, and pays $\nicefrac{{1}}{{3}}+\epsilon$ when the outcome $(D,C)$ is obtained. Effectively, player $1$ blocks player $2$ from incentivizing player $1$’s agent to cooperate, so player $2$’s best response is not to pay anything and get $\nicefrac{{1}}{{3}}+\epsilon$. This is an $\epsilon$-equilibrium for all $\epsilon>0$, since player $1$ incentivizes outcome $(D,C)$ in the limit $T\rightarrow\infty$, and he cannot improve his long-term payoff by more than $\epsilon$. The other equilibrium, as mentioned, is the mirror image of the same equilibrium (replacing the payment policies of players $1$ and $2$).
@@
Appendix C Additional Definitions
For completeness, we provide here several standard definitions.
Regret minimization:
For a given sequence of action profiles $s^{1},\dots,s^{T}$, the regret of agent $i$ is the difference in utility for $i$ between the actual utility in that sequence and the utility of the best fixed action in hindsight: $\text{Regret}_{i}(s^{1},\dots,s^{T})=\max_{s\in S_{i}}\big{[}\sum_{\tau=1}^{T}u_{i}(s,s_{-i}^{\tau})\big{]}-\sum_{\tau=1}^{T}u_{i}(s_{i}^{\tau},s_{-i}^{\tau}).$ A no-regret agent minimizes this quantity in the long term. There are several formulations for this property. We will use the following definition.
Definition 3.
An algorithm satisfies the (external) regret-minimization property if, for a time horizon parameter $T$ and any $T$-sequence of play of the other players $(s_{-i}^{1},\dots,s_{-i}^{T})$, where $s_{i}^{t}$ and $s_{-i}^{t}$ denote the actions taken at time $t$ by the algorithm and by the other players, respectively, we have that $\max_{s\in S_{i}}\big{[}\sum_{\tau=1}^{T}u_{i}(s,s_{-i}^{\tau})\big{]}-\sum_{\tau=1}^{T}u_{i}(s_{i}^{\tau},s_{-i}^{\tau})=o(T)$ with probability 1 in the limit as $T\rightarrow\infty$. An agent is regret-minimizing if it satisfies the regret-minimization property.
Coarse correlated equilibrium:
Coarse correlated equilibria are a weaker notion than correlated equilibria [5], also known as the Hannan set or Hannan consistent distributions [43]. The simplest definition for our purpose is the following.
Definition 4.
A joint distribution of play is a coarse correlated equilibrium if, under this distribution, all the players have in expectation regret at most zero.
Mean-based learning algorithms:
Mean-based learning algorithms [19] are a family of algorithms that play actions that are best responses to the history of play with high probability. This family was shown to include many standard no-regret algorithms, like multiplicative weights (see [3] and references therein), follow the perturbed leader [43, 51], and EXP3 [4].
Definition 5.
[19]. Let $\sigma_{i,t}=\sum_{t^{\prime}=1}^{t}u_{i,t^{\prime}}$, where $u_{i,t}$ is the utility of action $i$ at time $t$. An algorithm for the experts problem or the multi-armed bandits problem is $\gamma(T)$-mean-based if it is the case that whenever $\sigma_{i,t}<\sigma_{j,t}-\gamma(T)T$, then the probability that the algorithm pulls arm $i$ in round $t$ is at most $\gamma(T)$. An algorithm is mean-based if it is $\gamma(T)$-mean-based for some $\gamma(T)=o(1)$.
Appendix D Proofs for Section 3
D.1 Second-price auctions
Proof.
(Theorem 1). Consider a repeated second price auction with a single item sold in every step, and $n$ players with values $v_{1}>v_{2}\geq\dots\geq v_{n}$. Assume that the total payment that a player makes is bounded by some large constant $M>2nv_{1}$. Consider the following profile of payment policies. Every player $j>1$ uses the policy of always paying zero. Choose $\epsilon$ with $0<\epsilon<(v_{1}-v_{2})/n$. The payment policy for player $1$ is specified by the following conditions. (1) Whenever $b_{1}=v_{1}$, make a payment $p_{1j}$ to every player $j>1$ of $\epsilon/(n-1)$ if $b_{j}=0$ and zero otherwise. (2) If agent $1$ bids $b_{1}\neq v_{1}$, then he pays $p_{1j}=\frac{M}{n-1}$ to every other player $j>1$. Note that the maximum payment according to these conditions is $M$, as required.
Claim 1.
Proof.
Claim 2.
Proof.
We begin by looking at auctions with $n>2$. In every deviation by some agent $j>1$, in order to ensure that a bid $b<v_{1}$ is not dominated for agent $1$, agent $j$ must pay agent $1$ an amount of at least $M-v_{1}$. The utility for player $j$ in this case is bounded from above by $v_{j}-(M-v_{1})+\frac{M}{n-1}<2v_{1}-\frac{n-2}{n-1}M<2v_{1}\big{(}1-\frac{n(n-2)}{n-1}\big{)}<0$. That is, in every deviation by agent $j$ in which $j$ can win and get a positive utility from the auction (by incentivizing agent $1$ to bid below $v_{j}$ with some probability), the payment is too large, so the resulting utility is negative. So player $j$ prefers not to make payments and get a utility of $\epsilon/(n-1)$.
To see this, notice that a bid for which player $2$ pays agent $1$ an amount less than $M$ is still dominated and therefore not played by agent $1$. Suppose that there is a bid $b$ for which agent $2$ pays agent $1$ an amount of $M$. When agent $1$ loses with the bid $b$, he gets a utility of zero. When agent $1$ wins, he has an expected payoff of $v_{1}-\operatorname{\mathbb{E}}[b_{2}|b>b_{2}]$. Denote by $q$ the probability that $b$ is a winning bid for agent $1$. We have $v_{1}-\epsilon\leq q(v_{1}-\operatorname{\mathbb{E}}[b_{2}|b>b_{2}])\leq qv_{1},$ so we have $q\geq 1-\frac{\epsilon}{v_{1}}.$ Without loss of generality, we set $v_{1}=1$, and so agent $1$ wins at least $1-\epsilon$ fraction of the time.
Now, consider agent $2$. By the argument above, agent $2$’s payoff from using bids that are above zero is at most $v_{2}\cdot\epsilon<\epsilon$. Therefore, using such bids gives agent $2$ positive regret compared to bidding zero (when he bids zero, he gets a payment of $\epsilon$). Thus, agent $2$ almost always bids zero. It follows that player $2$ cannot make payments that increase his utility, so the policy of always paying zero is a best response. ∎
We are now almost done. With the policy profile that we have, player $1$ gets a utility of $v_{1}-\epsilon$. By taking a sufficiently small value of $\epsilon$, player $1$ gets a utility per time step that is arbitrarily close to $v_{1}$. Since $v_{1}$ is the optimal welfare in the game and player $1$ gets $\epsilon$ close to it for all $\epsilon>0$, this is an approximate best response for player $1$, and we have an $\epsilon$-Nash equilibrium. The auctioneer’s revenue in this equilibrium is zero, other than vanishing profits during the initial learning phase of the agents. ∎
D.2 First-price auctions
Proof.
(Lemma 1). We start by assuming that in the game between these agents, parameterized by $\eta$, there is a mixed Nash equilibrium with continuous CDFs, PDFs with full support on $[0,v_{2}-\eta]$, and player $2$ has a positive probability of bidding zero.
Consider player $1$’s utility. In equilibrium, he is indifferent between bidding zero and bidding $v_{2}-\eta$, giving us the size of the point mass for player $2$ at zero:
[…]
Player $1$’s utility when bidding $x$ is $u_{1}(x,y)=G_{2}(x)(v_{1}-x)-\chi(y)\eta$, where $\chi(y)$ is an indicator which equals $1$ if player $2$ bids zero (i.e., if $y=0$) and equals zero otherwise. Importantly, it is independent of $x$. Player $1$ is indifferent between all his bids, so $\frac{\partial u_{1}(x,y)}{\partial x}=0$, which gives
$G^{\prime}_{2}(x)(v_{1}-x)-G_{2}(x)=0.$
$c=v_{2}G_{2}(0)=v_{1}-v_{2}+\eta.$
Checking the consistency of the support (the CDF must be equal to $1$ at the top of the support):
$G_{2}(v_{2}-\eta)=\frac{v_{1}-v_{2}+\eta}{v_{1}-(v_{2}-\eta)}=1.$
Thus, we have a mixed Nash equilibrium with CDFs $F_{1}(x)=\frac{\eta}{v_{2}-x},\quad G_{2}(x)=\frac{v_{1}-v_{2}+\eta}{v_{1}-x}$, as stated in the lemma. ∎
Proof.
[…]
A straightforward calculation by differentiating this expression with respect to $\eta$ shows that the expectation is maximized when $\eta=v_{2}/2$, in which case its value is $v_{1}-v_{2}+v_{2}^{2}/4v_{1}$, as claimed. With this choice of $\eta$, player 2’s utility is $v_{1}/2$.
The frequency at which the low agent (agent $2$) wins is given by:
[…]
Proof.
Claim 3.
Proof.
Continuing the proof of Lemma 3, suppose by way of contradiction that $F_{1}(b^{\prime})>F_{1}^{NE}(b^{\prime})$ for some $b^{\prime}<v_{2}-\eta$. $F_{1}(b^{\prime})>F_{1}^{NE}(b^{\prime})$ implies $u_{2}(b^{\prime})=F_{1}(b^{\prime})(v_{2}-b^{\prime})>F_{1}^{NE}(b^{\prime})(v_{2}-b^{\prime})=\eta=u_{2}(0)$. Therefore, bidding zero is not in the support of player $2$.
In the standard first-price auction, without payments, mean-based agents can converge only to CCEs that include bids above $v_{2}-\eta$ in their support (see [53]), a contradiction. Therefore, we have $F_{1}=F_{1}^{NE}$.
Next, suppose by way of contradiction that $G_{2}(b^{\prime})>G_{2}^{NE}(b^{\prime})$ for some $b^{\prime}<v_{2}-\eta$. This implies that $u_{1}(b^{\prime})=G_{2}(b^{\prime})(v_{1}-b^{\prime})-\eta G_{2}(0)>G_{2}^{NE}(b^{\prime})(v_{1}-b^{\prime})-\eta G_{2}(0)=v_{1}-v_{2}+\eta-\eta G_{2}(0)=u_{1}(v_{2}-\eta)$. That is, bidding $b^{\prime}$ gives player 1 a higher utility than bidding $v_{2}-\eta$, so the latter is not in the support of player 1, which is a contradiction, since the marginal distribution of player $1$ is $F_{1}$, which is supported on $[0,v_{2}-\eta]$. Thus, overall, we have $F_{1}=F_{1}^{NE}$ and $G_{2}=G_{2}^{NE}$. ∎
Proof.
(Theorem 2.) We will use the fact that the game analyzed in our proof is induced by a unilateral payment policy by the high-value player. In any equilibrium, player $1$ can consider the deviation in which he rejects (pays back) all the payments he received and makes this unilateral payment, leading to a utility of $v_{1}-v_{2}+v_{2}^{2}/4v_{1}$. Therefore, in any equilibrium, player $1$ gets at least this utility, which is higher than the utility of $v_{1}-v_{2}$ that he gets in the game without payments. As for the utility of the low-value player, since the high-value player’s utility is above the second price, player $1$ must use positive payments, otherwise the dynamics of mean-based agents could approach only the second-price outcome (as shown in [53]). Therefore, in the equilibrium of the payment-policy game, the low-value player has positive utility, an improvement over the game without payments.
The fact that the revenue decreases in any equilibrium is now straightforward. To see this explicitly, suppose by way of contradiction that the revenue $R$ is at least the revenue $v_{2}$ of the game without payments. The total welfare (of the players and the auctioneer) is $w=u_{1}+R+u_{2}\leq v_{1}$. We then get $v_{1}+\eta\Big{(}\frac{v_{1}-\eta}{v_{1}}\Big{)}+u_{2}\leq v_{1}$, which is a contradiction, since player $2$ cannot get negative utility in equilibrium.
D.2.1 Extension of Lemma 3 to $n$ players.
@@ D.2 First-price auctions
To extend the lemmas in the proof of Theorem 2 to the case of $n$ players, consider the following analysis. Suppose we have $n$ players with values $v_{1}>v_{2}>v_{3}>\dots>v_{n}$, each using a regret-minimizing agent. If agent $2$ bids $v_{3}$, agent $1$ pays him $\eta>0$.
@@
As in the proof of Lemma 1, we begin by assuming that in the game between these agents, parameterized by $\eta$, there is a mixed Nash equilibrium with continuous CDFs, PDFs with full support on $[v_{3},v_{2}-\eta]$, and player $2$ has a positive probability of bidding $v_{3}$.
We start by looking at player $1$’s utility. In equilibrium, he is indifferent between bidding $v_{3}$ and bidding $v_{2}-\eta$, giving us the (point-mass) probability that player $2$ bids $v_{3}$:
@@
Player $1$’s utility when bidding $x$ is $u_{1}(x,y)=G_{2}(x)(v_{1}-x)-\chi(y)\eta$, where $\chi(y)$ is an indicator which equals $1$ if player $2$ bids zero and equals zero otherwise. As before, note that it is independent of $x$. Player $1$ is indifferent between all his bids, so $\frac{\partial u_{1}(x,y)}{\partial x}=0$, which gives
$G^{\prime}_{2}(x)(v_{1}-x)-G_{2}(x)=0.$
We have an ordinary differential equation with a unique solution of the form
@@
So we see that in this equilibrium, player $1$ manages to increase his payoff with a unilateral payment. The optimal payment is $\eta^{*}=\min(v_{2}/2,v_{2}-v_{3})$. The payoff for player $2$ is $\eta$. Given the analysis above of Lemmas 1 and 2 with $n$ agents, the extension of Lemma 3 holds without any change to the proof, noticing the fact that since the minimum of the support is $v_{3}$, any player $j>2$ never wins the auction, so these agents can have arbitrary distributions with support at most $v_{3}$ with mean-based dynamics. Since this is a utility-improving deviation for player $1$ from the zero-utility payment profile, the auction is not stable with $n$ players.
Appendix E Proofs for Section 4
Proof.
(Lemma 4). We start constructing a payment policy for the agent of player $i$ by looking at the optimal $k$-implementation payments for inducing $y$ as a dominant strategy equilibrium; we will then adjust this payment profile to fit our setting in which there are no external payments and to minimize player $i$’s cost. A key point is that the payment amount $k(y)$ includes payments to agent $i$ that an external designer would have made for inducing $y_{i}$ as the dominant strategy for agent $i$. These payments are not feasible and not needed in our case. Instead, we change the payment policy of agent $i$ in the following way: for any action $s_{i}\neq y_{i}$, agent $i$ makes a payment of $\max_{s\in S}u_{i}(s)+1$, and distributes this payment equally to the other agents. These payments make all the other actions of agent $i$ strictly dominated by the action $y_{i}$.
Thus, to calculate the cost for player $i$, we need to subtract from the $k$-implementation payments the parts that are paid to player $i$, $R_{i}(y,y)$, and to add a learning-phase error term $\tilde{c}$ that results from the additional payments that agent $i$ makes when playing its dominated strategies, as well as from other times during the learning dynamics where the other agents play their dominated strategies.
Claim.
The expected learning-phase cost-per-time-step $\tilde{c}$ vanishes in the long term, that is, $\tilde{c}=o(1)$.
Proof.
$\tilde{c}$ is the expected total payment per time step that player $i$ makes when agents play strictly dominated strategies. Assume by way of contradiction that $\lim_{T\rightarrow\infty}\tilde{c}>0$, or that the limit does not exist. Both assumptions imply that the frequency at which dominated strategies are played in the agents’ dynamic is not vanishing, which contradicts the regret-minimization property of the agents. ∎
Proof.
(Lemma 5). We need to specify the requirement that player $i$ will increase his own utility after making the payments that induce $y$ as the unique unique outcome of the dynamics. That is, $cost_{i}(y)<u_{i}(y)-u_{i}(x)$. By Lemma 4 we have that for large enough $T$, player $i$ prefers making such payments if $u_{i}(y)-k(y)+R_{i}(y,y)>u_{i}(x)$. Rewriting the left terms we have
$u_{i}^{BR}(y)-k(y)>u_{i}(x)=w(x)-\sum_{j\neq i}u_{j}(x).$
Using the fact that, by definition [62], $k(y)=\big{(}\sum_{j}u_{j}^{BR}(y)\big{)}-w(y)$, we can rewrite the left-hand side as $w(y)-\sum_{j\neq i}u_{j}^{BR}(y).$ Rearranging the terms, we have
Proof.
Suppose that $y^{*}$ is a Nash equilibrium. By the definition of Nash equilibrium, if some profile $s$ is a Nash equilibrium, then $R_{i}(s,s)=0$ for all $i$. Thus, by Lemma 4, any player can construct a payment policy that induces $y^{*}$ as the unique outcome of the agents’ dynamic with a cost-per-time-step of $o(1)$.
Therefore, to have a player $i$ who increases his payoff by inducing $y^{*}$ as the outcome of the dynamics, we only need to have that $u_{i}(y^{*})>u_{i}(x)$. Assume by way of contradiction that for all $i$, $u_{i}(x)\geq u_{i}(y^{*})$. Summing over the players, we have $\sum_{i}u_{i}(x)\geq\sum_{i}u_{i}(y^{*})$; that is, $w(x)\geq w(y^{*})$, a contradiction to the optimality of $y^{*}$. Thus, we have that for $T$ large enough that the learning-phase costs are small, there exists a player $i$ who can increase his payoff by inducing the optimal-welfare outcome as the outcome of the agents’ dynamics. If $y^{*}$ is a non-equilibrium outcome, the conditions and proof are given by Lemma 5 with $y=y^{*}$. Notice also that since under this payment policy $y^{*}$ dominates $x$ for all the agents, the utility for all the players is higher than in the game without payments. ∎
Proof.
(Theorem 4). The result follows from combining (the proof of) Lemma 5 with general properties of regret-minimization dynamics. Let $\Gamma$ be a game in which the Nash equilibrium with the highest welfare, denoted $\tilde{y}$, yields a welfare of $\tilde{w}$ (where $\tilde{w}$ is not necessarily equal to the optimal welfare $OPT$), and there exists a CCE $x$ with welfare $w(x)<\tilde{w}$.
It is known that for any CCE of a game, there are regret-minimizing agents that converge to that equilibrium [54, 63]. In particular, there are agents that converge arbitrarily close to $x$ [^13] Technically, the proofs of this convergence result use cycles of pure actions to approximate the time average of joint distributions, and the arbitrary approximation is due to the density of the rational numbers. We disregard this detail in our analysis. By Lemma 5 (in particular, Corollary 1), there exists a player who prefers using a payment policy that assures convergence to the highest welfare equilibrium $\tilde{y}$ compared to not making payments and reaching outcome $x$. Thus, we have a utility-improving deviation over the no-payments action profile and hence the game is not stable. ∎
Appendix F Proofs for Section 5
Proof.
(Theorem 5). The intuition for the proof in the two-player case is that to have a positive payment in equilibrium, some player must benefit from making the payment, while the other benefits from receiving the payment. With more than two players, there might be payments that are mutually beneficial to some subset of players but have negative externalities for others. In particular, with more than two players, there might be players who cannot profitably counteract payments made among others that do not benefit them.
Formally, consider a game $\Gamma$, a set $A$ of regret-minimizing learning agents for the players, and a time horizon $T$. Suppose, for the sake of contradiction, that there exists an equilibrium $p^{*}$ of the payment-policy game associated with ${\Gamma,A,T}$ where player $i$ has a lower utility than their expected utility over the $T$ rounds in the game without payments. Let $h$ denote the utility for player $i$ in the game without payments, and let $l$ denote the utility for player $i$ in the equilibrium $p^{*}$. By our assumption, we have $h>l$. Consider the following deviation from the payment policy $p^{*}$: player $i$ matches his payments to zero out all payments made by the other player and, additionally, reduces all of his other payments to zero. As a result, the two agents now observe in their dynamics the utilities of the original game $\Gamma$ without payments. Therefore, player $i$ has a utility of $h$. Since player $i$ has increased his utility from $l$ to $h$, we have a contradiction to $p^{*}$ being an equilibrium. Hence, the utilities of both players can only increase relative to the game without payments. ∎
Proof.
(Theorem 6). This result is related to a theorem in [54] that deals with games with dominant strategies in a meta-game in which users misreport their utility parameters to their learning agents. In our context of payments between the agents, the result is more general, applying to a larger family of games, including all dominance-solvable games, generic fully-mixed games, and socially-concave games [34] (a sub-class of concave games in which regret-minimization dynamics converge to Nash equilibria) with a unique equilibrium.
Consider as our starting point the action profile where the players use zero payments. In the game without payments, any regret-minimization dynamic will converge to the unique coarse correlated equilibrium (which is also the unique Nash equilibrium). To show that the game is not stable for any set of regret-minimizing agents for the players, we need to find a utility-improving deviation from the zero-payment profile for one of the two players.
We will use the fact that when the opponent does not make any payments, a player has the ability to alter his policy so as to commit to playing a pure action in the long term; this can be obtained by making large payments whenever the agent of this player plays different actions, inducing a dominant strategy for the agent. Since for at least one of the players there is a Stackelberg-equilibrium outcome that is different from the (unique) Nash-equilibrium outcome, there exists a player who obtains a higher payoff as a Stackelberg leader in the game, so this player prefers making such a commitment. Since the other agent is also minimizing regret, this agent will eventually learn to best reply to the fixed action of the first player’s agent, and the dynamics with these payments will converge to the Stackelberg outcome of the game. Thus, we have a utility-improving deviation from the zero-payment policy, so the game is not stable. Note that this is a different result from that of Theorem 4, since the Stackelberg outcome here is not a coarse correlated equilibrium of the game $\Gamma$. ∎
Acknowledgments
We thank Noam Nisan for his valuable feedback. Éva Tardos was supported in part by AFOSR grant FA9550-23-1-0410, AFOSR grant FA9550-231-0068. Joe Halpern was supported in part by AFOSR grant FA23862114029, MURI grant W911NF-19-1-0217, ARO grant W911NF-22-1-0061, NSF grant FMitF-2319186, and a grant from the Cooperative AI Foundation.

### v2

@@ Appendix A Prisoner’s Dilemma Games
There are various ways to show this result. Consider the following induction argument. The game has at least one order of iterated elimination of strictly dominated strategies. Index the players according to one such order, such that player $1$ has a dominant strategy, player $2$ has a dominant strategy in the sub-game where player $1$ plays her dominant strategy, and so on. Additionally, index the corresponding dominant strategies of the players using the same order, so that $a_{1}$ is the first action that is taken, $a_{2}$ is the second, and so on. Clearly, in any CCE, player $1$ plays her dominant strategy $a_{1}$; otherwise she has positive regret. Our induction step is simple: Fix any CCE. Given that players $11,...,k-$ 1 play actions $a_{1},...,a_{k-1}$, player $k$ has a strict best response to play $a_{k}$. By induction, this is true up to the last player $n$,so we have that our arbitrary CCE is the pure Nash equilibrium profile $(a_{1},...,a_{n})$. Therefore, regret-minimization dynamics converge to it with high probability. ∎
@@
To approach the outcome described in the introduction as an equilibrium, we need to adjust the payment policy of player $1$ as follows. Player $1$’s agent pays player $2$’s agent the maximum payment whenever player $1$’s agent cooperates, and pays $\nicefrac{{1}}{{3}}+\epsilon$ when the outcome $(D,C)$ is obtained. Effectively, player $1$ blocks player $2$ from incentivizing player $1$’s agent to cooperate, so player $2$’s best response is not to pay anything and get $\nicefrac{{1}}{{3}}+\epsilon$. This is an $\epsilon$-equilibrium for all $\epsilon>0$, since player $1$ incentivizes outcome $(D,C)$ in the limit $T\rightarrow\infty$, and she cannot improve her long-term payoff by more than $\epsilon$. The other equilibrium, as mentioned, is the mirror image of the same equilibrium (replacing the payment policies of players $1$ and $2$).
@@
Appendix B Extension of Lemma 3
@@
To extend the lemmas of Theorem 2 to the case of $n$ players, consider the following analysis. Suppose we have $n$ players with values $v_{1}>v_{2}>v_{3}>\dots>v_{n}$, each using a regret-minimizing agent. If agent $2$ bids $v_{3}$, agent $1$ pays her $\eta>0$. Lemma 3 and 2 have the following simple adjustments.
@@
As in the proof of Lemma 1, we begin by assuming that in the game between these agents, parameterized by $\eta$, there is a mixed Nash equilibrium with continuous CDFs, PDFs with full support on $[v_{3},v_{2}-\eta]$, and player $2$ has a positive probability of bidding $v_{3}$. We start by looking at player $1$’s utility. In equilibrium, she is indifferent between bidding $v_{3}$ and bidding $v_{2}-\eta$, giving us the (point-mass) probability that player $2$ bids $v_{3}$:
@@
Player $1$’s utility when bidding $x$ is $u_{1}(x,y)=G_{2}(x)(v_{1}-x)-\chi(y)\eta$, where $\chi(y)$ is an indicator which equals $1$ if player $2$ bids zero and equals zero otherwise. As before, note that it is independent of $x$. Player $1$ is indifferent between all her bids, so $\frac{\partial u_{1}(x,y)}{\partial x}=0$, which gives $G^{\prime}_{2}(x)(v_{1}-x)-G_{2}(x)=0$. We have an ordinary differential equation with a unique solution of the form
@@
So we see that in this equilibrium, player $1$ manages to increase her payoff with a unilateral payment. The optimal payment is $\eta^{*}=\min(v_{2}/2,v_{2}-v_{3})$. The payoff for player $2$ is $\eta$.
@@
Given the analysis above of Lemmas 1 and 2 with $n$ agents, the extension of Lemma 3 holds without any change to the proof, noticing the fact that since the minimum of the support is $v_{3}$, any player $j>2$ never wins the auction, so these agents can have arbitrary distributions with support at most $v_{3}$ with mean-based dynamics. Since this is a utility-improving deviation for player $1$ from the zero-utility payment profile, the auction is not stable with $n$ players.
Appendix C Additional Definitions
For completeness, we provide here several standard definitions.
Regret minimization: For a given sequence of action profiles $s^{1},\dots,s^{T}$, the regret of agent $i$ is the difference in utility for $i$ between the actual utility in that sequence and the utility of the best fixed action in hindsight: $\text{Regret}_{i}(s^{1},\dots,s^{T})=\max_{s\in S_{i}}\sum_{\tau=1}^{T}u_{i}(s,s_{-i}^{\tau})-\sum_{\tau=1}^{T}u_{i}(s_{i}^{\tau},s_{-i}^{\tau}).$ A no-regret agent minimizes this quantity in the long term. There are several formulations for this property, including high probability definitions; we give the following definition for completeness.
Definition 3.
An algorithm satisfies the (external) regret-minimization property if, for a time horizon parameter $T$ and any $T$-sequence of play of the other players $(s_{-i}^{1},\dots,s_{-i}^{T})$, where $s_{i}^{t}$ and $s_{-i}^{t}$ denote the actions taken at time $t$ by the algorithm and by the other players, respectively, we have that in expectation over the actions taken by the algorithm, $\max_{s\in S_{i}}\sum_{\tau=1}^{T}u_{i}(s,s_{-i}^{\tau})-\sum_{\tau=1}^{T}u_{i}(s_{i}^{\tau},s_{-i}^{\tau})=o(T)$ as $T\rightarrow\infty$. An agent is regret-minimizing if it satisfies the regret-minimization property.
Coarse correlated equilibria: Coarse correlated equilibria are a weaker notion than correlated equilibria [8], also known as the Hannan set or Hannan consistent distributions [52]. See also [93]. The simplest definition for our purpose is the following.
Definition 4.
A joint distribution of play is a coarse correlated equilibrium if under this distribution all players have in expectation regret at most zero.
Mean-based learning algorithms: Mean-based learning algorithms [22] are a family of algorithms that play with high probability actions that are best responses to the history of play. This class was shown to include many standard no-regret algorithms, like multiplicative weights (see [5] and references therein), follow the perturbed leader [52, 61], and EXP3 [7].
Definition 5.
(From [22]). Let $\sigma_{i,t}=\sum_{t^{\prime}=1}^{t}u_{i,t^{\prime}}$, where $u_{i,t}$ is the utility of action $i$ at time $t$. An algorithm for the experts problem or the multi-armed bandits problem is $\gamma(T)$-mean-based if it is the case that whenever $\sigma_{i,t}<\sigma_{j,t}-\gamma(T)T$, then the probability that the algorithm pulls arm $i$ in round $t$ is at most $\gamma(T)$. An algorithm is mean-based if it is $\gamma(T)$-mean-based for some $\gamma(T)=o(1)$.